The area in square meters for a triangular
sail is given by A(x) =x^2 + 5x + 6.
Find A(5).
b) If the height of the sail is x +3 meters, then what is
the length of the base of the sail?

Answer by RAY100(1637)

A(x) = x^2 +5x +6
.
A(5) = 5^2 +5*5 +6 = 56
.
A = (1/2)bh = (1/2) b (x+3)
A(5) = 56 = (1/2) 8 b = 4b
b= 56/4=14
.
check A = .5 * 14 * 8 = 56,,,ok